# Enabling Uncertainty Estimation in Iterative Neural Networks

### v1

@@ 1 Introduction
Figure 1: Uncertainty in recursive models. Such models use their initial predictions as inputs to produce subsequent predictions. We display the output of three consecutive iterations of a network trained to compute distance maps to road pixels. (Top:) All roads are clearly visible. The three maps are similar and the per pixel variance is low. (Bottom:) The road in the red square is tree-covered. It is eventually detected properly but the variance is high.
It has long been known that using deep networks to recursively refine predictions is often beneficial. This has been demonstrated for semantic segmentation (Zhou et al., 2018; Wang et al., 2019), pose estimation (Newell et al., 2016), depth estimation (Zhang et al., 2018a), multi-task learning (Xu et al., 2018), delineation (Mosińska et al., 2018), natural language processing (A.Vaswani et al., 2017; Devlin et al., 2018), among others. Given a network $f$ parameterized by weights $\Theta$ and takes as input a vector $\mathbf{x}$, which can represent an image or a text, and produces an output $\mathbf{y}$, the output of the recursion’s $i^{\rm th}$ iteration can be written as $\mathbf{y}_{i}=f_{\Theta}(\mathbf{x},\mathbf{y}_{i-1})$, where $\mathbf{y}_{i-1}$ is the output of the previous iteration. This recursion often yields improved performance over non-iterative methods using the same network architectures (Shen et al., 2017; Zhang et al., 2018b; Wang et al., 2019; Oner et al., 2021, 2022), while requiring less labeled data for training purposes (Mosińska et al., 2018). In essence, giving a previous output as input to the network sets up a virtuous circle in which the model receives relevant spatial attention signals that serve as priors and help generate improved predictions. For example, in the road delineation example of Fig. 1, the presence of road fragments with gaps in them cues the network to the possible existence of connecting segments.
@@
More specifically, we derive uncertainty measures by analyzing the variance in outputs from consecutive iterations of an iterative model, where higher variance indicates greater uncertainty. We demonstrate that this is fast, accurate, and easy to deploy in two very different application domains, road detection in aerial images and the estimation of aerodynamic properties of 2D and 3D shapes. These two applications feature unique sets of challenges, which underscore the versatility and broad applicability of our approach.
@@ 2.1 Uncertainty Estimation
Deep Ensembles involve training multiple networks, starting from different initial conditions. They are noted for the potential diversity of their predictions, attributable to randomness from weight initialization, data augmentation, and stochastic gradient updates. This diversity is central to their effectiveness (Perrone & Cooper, 1995; Fort et al., 2019). In many situations, they deliver more reliable uncertainty estimates than other methods (Ovadia et al., 2019; Gustafsson et al., 2020; Ashukha et al., 2020; Postels et al., 2022). They therefore remain the leading technique, despite the high computational cost of training several networks instead of a single one and of performing several forward passes at inference time.
Among the other techniques, MC-Dropout involves randomly zeroing out network weights and assessing the effect and is popular due to its lower computational cost. Unfortunately, its estimates remain less reliable than those of Deep Ensembles (Ashukha et al., 2020), even though there has been recent attempts at improving it (Wen et al., 2020; Durasov et al., 2021). Similarly, Bayesian Networks rarely outperform Deep Ensembles (Blundell et al., 2015; Graves, 2011; Hernández-Lobato & Adams, 2015; Kingma et al., 2015).
All the above methods are sampling-based and require several forwards passes at inference time. Thus, when a fast response time is required, as in robotics control, sampling-free approaches with single-pass inference become of interest. However, deploying them often requires significant modifications to the network’s architecture (Postels et al., 2019), substantial changes to the training procedures (Malinin & Gales, 2018), limiting their application to very specific tasks (Amersfoort et al., 2020; Malinin & Gales, 2018; Mukhoti et al., 2021), or reducing the quality of the uncertainty estimate (Postels et al., 2022; Ashukha et al., 2020). As a result, they have not gained as much traction as MC-Dropout and Ensembles.
@@ 2.2 Iterative Refinement Methods
Iterative refinement techniques have been for many different purposes (Mnih & Hinton, 2010; Pinheiro & Collobert, 2014; Tu & Bai, 2009; Shen et al., 2017). This incorporates surrounding context into the prediction (Seyedhosseini et al., 2013), proving particularly useful for tasks such delineation  (Sironi et al., 2016), human pose estimation (Newell et al., 2016), semantic segmentation (Zhang et al., 2018b; Wang et al., 2019), depth estimation (Durasov et al., 2019; Xu et al., 2018), and multi-task learning (Durasov et al., 2022b). In the first use case of this work, we build upon the recursive networks used in (Mosińska et al., 2018) to delineate roads by computing distance maps to the road pixels, as shown in Fig. 1. The network is a UNet (Ronneberger et al., 2015) that takes as input the image and the distance map computed at the previous iteration, starting from a blank one. However, whereas the typical focus of delineation papers is to increase performance in terms of a number of delineation metrics, ours is to provide an uncertainty estimate on the detections without sacrificing performance. In the second use case, we use an iterative model with Graph Neural Networks (Monti et al., 2017; Baqué et al., 2018) for 2D and 3D shape optimization. We show that uncertainty measures can be used to effectively select out-of-distribution data to enhance the training dataset.
@@ 3.1 Motivation
  • In distribution, with aleatoric noise. $\mathbf{x}$ is in-distribution, the training data is noisy, making the he aleatoric uncertainty higher and the probability $p$ less peaked. If $\mathbf{y}_{i}$ is not initially correct but still within the main mode of the probability distribution, the derivative will be smaller than in the no-noise scenario, and thus the convergence slower, as in the case of $x=5.0$.
  • Out of distribution. $\mathbf{x}$ is out-of-distribution, which can be understood as epistemic uncertainty and $p$ cannot be expected to have a well-defined peaked shape but often tends to be flatter. The behavior is then somewhat random as in the case of $x=-2.0$ and $x=7.0$.
@@ 3.2 Estimating Uncertainty
In the result section, we show that this variance estimate strongly correlates with the actual accuracy of the prediction on experimental data. Note that for the arguments made for two in-distribution cases discussed at the end Section 3.1 to apply, the prediction $\mathbf{y}_{i}$ has to fall within the mode of the probability distribution. To maximize the chances of this happening, at training time, we supervise the network so that all $\mathbf{y}_{i}$ in $\mathbf{Y}$ are as close possible to the ground-truth by minimizing
@@
where $D$ is a measure of distance and gt stands for ground-truth.
@@ 4.1 Delineation
For road delineation, the true measure of success is preservation of the topology of the road networks rather than the very precise location of the centerline. This is ussually expressed in terms of hte following metrics: APLS ($\uparrow$). Average Path Length Similarity, defined as an aggregation of relative length difference of shortest paths between pairs of corresponding points in the reconstructed and predicted maps (van Etten, 2019). CCQ ($\uparrow$). Metric that measures spatial co-occurrence of annotated and predicted road pixels. The Correctness, Completeness and Quality are equivalent to precision, recall and intersection-over-union, where the definition of a true positive has been relaxed from spatial coincidence of prediction and annotation to co-occurrence within a distance of 5 pixels (Wiedemann et al., 1998). F1 Score ($\uparrow$). A balance between precision and recall, the F1 score is twice the product of precision and recall divided by their sum. It’s widely used in binary segmentation to equally weigh false positives and false negatives (Fawcett, 2006).
@@ 4.2 Aerodynamics Prediction and Optimization
We use the following metrics to evaluate the quality of our baselinse in terms of predictions accuracy and uncertainty estimation: Mean absolute error ($\downarrow$) (MAE). It is the average of the absolute differences between the predicted and actual values. It is a common metric for regression tasks. Opimized performance. We use our baselines to perform Bayesian optimization as it was described above. We then report the best performance value, $\textbf{r}_{i}$ lift-to-drag value for airfoils and $\textbf{r}_{i}$ drag value for cars, obtained by each method and the dynamics of optimization process.
@@
Table 4: Evaluation the uncertainty measure for 2D airfoils (AIR) and 3D cars (CAR). The best result in each category is in bold and the second best is in bold. They all correspond to DEBOSH and Deep Ensembles. The two approaches are comparable in terms of evaluating uncertainty but the reentrant GNNs deliver better accuracy, as shown in Fig. 5.
@@ A.4 Propagating Information
Figure 10: Propagating information across an shape. A comparable behavior is observed when increasing the depth of a standard GNN (red curves) and when running several iterations of a shallower iterative GNN (black curves).

### v2

@@ 1 Introduction
Figure 1: Uncertainty in recursive models. Such models use their initial predictions as inputs to produce subsequent predictions. We display the output of three consecutive iterations of a network trained to compute distance maps to road pixels. (Top:) All roads are clearly visible. The three maps are similar and the per-pixel variance is low. (Bottom:) The road in the red square is tree-covered. It is eventually detected properly but the variance is high.
It has long been known that using deep networks to recursively refine predictions is often beneficial. This has been demonstrated for semantic segmentation (Zhou et al., 2018; Wang et al., 2019), pose estimation (Newell et al., 2016), depth estimation (Zhang et al., 2018a), multi-task learning (Xu et al., 2018), delineation (Mosińska et al., 2018), natural language processing (A.Vaswani et al., 2017; Devlin et al., 2018), among others. Given a network $f$ parameterized by weights $\Theta$ and takes as input a vector $\mathbf{x}$, which can represent an image or a text, and produces an output $\mathbf{y}$, the output of the recursion’s $i^{\rm th}$ iteration can be written as $\mathbf{y}_{i}=f_{\Theta}(\mathbf{x},\mathbf{y}_{i-1})$, where $\mathbf{y}_{i-1}$ is the output of the previous iteration. This recursion often yields improved performance over non-iterative methods using the same network architectures (Shen et al., 2017; Zhang et al., 2018b; Wang et al., 2019; Oner et al., 2021, 2022), while requiring less labeled data for training purposes (Mosińska et al., 2018). In essence, giving a previous output as input to the network sets up a virtuous circle in which the model receives relevant spatial attention signals that serve as priors and help generate improved predictions. For example, in the road delineation example of Fig. 1, the presence of road fragments with gaps in them cues the network to the possible existence of connecting segments. These ideas have been explored before the advent of Deep Learning, for example using Tensor Voting (Medioni et al., 2000), but incorporating them into deep networks has given them a new lease on life.
@@
More specifically, we derive uncertainty measures by analyzing the variance in outputs from consecutive iterations of an iterative model, where higher variance indicates greater uncertainty. We demonstrate that this is fast, accurate, and easy to deploy in two very different application domains, road detection in aerial images and the estimation of aerodynamic properties of 2D and 3D shapes. These two applications feature unique sets of challenges, which underscore the versatility and broad applicability of our approach. computational demands compared to Ensembles.
@@ 2.1 Uncertainty Estimation
Deep Ensembles involve training multiple networks, starting from different initial conditions. They are noted for the potential diversity of their predictions, attributable to randomness from weight initialization, data augmentation, and stochastic gradient updates. This diversity is central to their effectiveness (Perrone & Cooper, 1995; Fort et al., 2019). In many situations, they deliver more reliable uncertainty estimates than other methods (Ovadia et al., 2019; Gustafsson et al., 2020; Ashukha et al., 2020; Postels et al., 2022). They, therefore, remain the leading technique, despite the high computational cost of training several networks instead of a single one and of performing several forward passes at inference time. One of the active research directions is reducing the training and inference time of ensembling methods, as well as their memory requirements. For example, Antorán et al. (2020) try to emulate ensemble predictions by producing several outputs based on features at different levels of the model, and Daxberger et al. (2021) perform Bayesian inference only on a subset of the model’s weights chosen through a pruning-like procedure.
Among the other techniques, MC-Dropout involves randomly zeroing out network weights and assessing the effect and is popular due to its lower computational cost. Unfortunately, its estimates remain less reliable than those of Deep Ensembles (Ashukha et al., 2020), even though there have been recent attempts at improving it (Wen et al., 2020; Durasov et al., 2021). Similarly, Bayesian Networks rarely outperform Deep Ensembles (Blundell et al., 2015; Graves, 2011; Hernández-Lobato & Adams, 2015; Kingma et al., 2015).
All the above methods are sampling-based and require several forward passes at inference time. Thus, when a fast response time is required, as in robotics control, sampling-free approaches with single-pass inference become of interest. However, deploying them often requires significant modifications to the network’s architecture (Postels et al., 2019), substantial changes to the training procedures (Malinin & Gales, 2018), limiting their application to very specific tasks (Amersfoort et al., 2020; Malinin & Gales, 2018; Mukhoti et al., 2021), or reducing the quality of the uncertainty estimate (Postels et al., 2022; Ashukha et al., 2020). As a result, they have not gained as much traction as MC-Dropout and Ensembles.
@@ 2.2 Iterative Refinement Methods
Iterative refinement techniques have been used for many different purposes (Mnih & Hinton, 2010; Pinheiro & Collobert, 2014; Tu & Bai, 2009; Shen et al., 2017). This incorporates surrounding context into the prediction (Seyedhosseini et al., 2013), proving particularly useful for tasks such delineation  (Sironi et al., 2016), human pose estimation (Newell et al., 2016), semantic segmentation (Zhang et al., 2018b; Wang et al., 2019), depth estimation (Durasov et al., 2019; Xu et al., 2018), and multi-task learning (Durasov et al., 2022b). In the first use case of this work, we build upon the recursive networks used in (Mosińska et al., 2018) to delineate roads by computing distance maps to the road pixels, as shown in Fig. 1. The network is a UNet (Ronneberger et al., 2015) that takes as input the image and the distance map computed at the previous iteration, starting from a blank one. However, whereas the typical focus of delineation papers is to increase performance in terms of a number of delineation metrics, ours is to provide an uncertainty estimate on the detections without sacrificing performance. In the second use case, we use an iterative model with Graph Neural Networks (Monti et al., 2017; Baqué et al., 2018) for 2D and 3D shape optimization. We show that uncertainty measures can be used to effectively select out-of-distribution data to enhance the training dataset.
@@ 3.1 Motivation
  • In distribution, with aleatoric noise. $\mathbf{x}$ is in-distribution, the training data is noisy, making aleatoric uncertainty higher and the probability $p$ less peaked. If $\mathbf{y}_{i}$ is not initially correct but still within the main mode of the probability distribution, the derivative will be smaller than in the no-noise scenario, and thus the convergence is slower, as in the case of $x=5.0$.
  • Out of distribution. $\mathbf{x}$ is out-of-distribution, which can be understood as epistemic uncertainty, and $p$ cannot be expected to have a well-defined peaked shape but often tends to be flatter. The behavior is then somewhat random as in the case of $x=-2.0$ and $x=7.0$.
@@ 3.2 Estimating Uncertainty
In the result section, we show that this variance estimate strongly correlates with the actual accuracy of the prediction on experimental data. Note that for the arguments made for two in-distribution cases discussed at the end of Section 3.1 to apply, the prediction $\mathbf{y}_{i}$ has to fall within the mode of the probability distribution. To maximize the chances of this happening, at training time, we supervise the network so that all $\mathbf{y}_{i}$ in $\mathbf{Y}$ are as close as possible to the ground truth by minimizing
@@
where $D$ is a measure of distance and gt stands for ground truth.
@@ 4.1 Delineation
For road delineation, the true measure of success is preservation of the topology of the road networks rather than the very precise location of the centerline. This is ussually expressed in terms of hte following metrics:
  • APLS ($\uparrow$). Average Path Length Similarity, defined as an aggregation of relative length difference of shortest paths between pairs of corresponding points in the reconstructed and predicted maps (van Etten, 2019).
  • CCQ ($\uparrow$). Metric that measures spatial co-occurrence of annotated and predicted road pixels. The Correctness, Completeness and Quality are equivalent to precision, recall and intersection-over-union, where the definition of a true positive has been relaxed from spatial coincidence of prediction and annotation to co-occurrence within a distance of 5 pixels (Wiedemann et al., 1998).
  • F1 Score ($\uparrow$). A balance between precision and recall, the F1 score is twice the product of precision and recall divided by their sum. It’s widely used in binary segmentation to equally weigh false positives and false negatives (Fawcett, 2006).
@@ 4.2 Aerodynamics Prediction and Optimization
We use the following metrics to evaluate the quality of our baselinse in terms of predictions accuracy and uncertainty estimation:
  • Mean absolute error ($\downarrow$) (MAE). It is the average of the absolute differences between the predicted and actual values. It is a common metric for regression tasks.
  • Opimized performance. We use our baselines to perform Bayesian optimization as it was described above. We then report the best performance value, $\textbf{r}_{i}$ lift-to-drag value for airfoils and $\textbf{r}_{i}$ drag value for cars, obtained by each method and the dynamics of optimization process.
@@
Table 4: Evaluation the uncertainty measure for 2D airfoils (AIR) and 3D cars (CAR). The best result in each category is in bold and the second best is in bold. They all correspond to Ours and Deep Ensembles. The two approaches are comparable in terms of evaluating uncertainty but the reentrant GNNs deliver better accuracy, as shown in Fig. 5.
@@
5 Ablation study
5.1 Calibration Evaluation
Though rAULC and correlation metrics from Sec. 4 provide comprehensive information about calibration quality, we also provide the results for Expected Calibration Error (ECE) (Guo et al., 2017) evaluation. Tab. 5 provides ECE values for different methods, including MC-Dropout, Ensembles, and our proposed method. The results show that our method achieves better calibration than both Ensembles and MC-Dropout for all datasets.
MC-Dropout Ensembles Ours
ECE (RT) 0.997 1.138 0.475
ECE (MS) 0.558 0.794 0.419
ECE (Airfoils) 1.758 1.162 1.142
ECE (Cars) 0.267 0.232 0.227
Table 5: Expected Calibration Error (ECE) evaluation. As it was previously demonstrated for the rAULC metric, our method outperforms other approaches in terms of ECE calibration.
5.2 Image Classification
We expanded our experimental evaluation to include the widely-used task of image classification on the MNIST dataset, a popular benchmark for out-of-distribution (OOD) detection and uncertainty estimation. For the OOD task, we used FashionMNIST as the OOD samples, which is a standard choice in this context. To add more variability, we ran these experiments with two model architectures:
  • CNN Architecture: A convolutional neural network (CNN) with several 2D convolutional layers followed by a fully-connected classification head.
  • MLP Architecture: A multilayer perceptron (MLP) with 5 fully-connected layers, treating images as 784-dimensional vectors.
We evaluated these experiments using the same metrics as in our previous evaluations, ensuring consistency and comparability. The results are summarized in Table 6. As for our previous results, our method outperforms other approaches both in terms of model’s accuracy and uncertainty quality.
MC-DP Ensembles Ours
Acc 97.8 98.8 99.0 CNN
ECE 0.021 0.018 0.013
ROC-AUC 94.4 98.5 98.5
PR-AUC 93.8 98.1 98.5
Train 1x 5x 3x
Inf 5x 5x 3x
Acc 96.1 97.7 97.4 MLP
ECE 0.026 0.022 0.020
ROC-AUC 61.5 89.5 89.6
PR-AUC 69.2 89.9 88.7
Train 1x 5x 3x
Inf 5x 5x 3x
Table 6: MNIST classification results for CNN (top), and MLP (bottom) architectures. The best result in each category is in bold and the second best is in bold. Most correspond to Ours and Deep Ensembles.
@@ A.4 Propagating Information
Figure 10: Propagating information across a shape. A comparable behavior is observed when increasing the depth of a standard GNN (red curves) and when running several iterations of a shallower iterative GNN (black curves).
@@ A.5 Gradient Optimization
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
A.6 Ensembles on Iterative Architecture
Another baseline for comparison could involve applying Deep Ensembles (DE) to our iterative architecture. While this approach would significantly increase computational complexity, it could provide better metrics for uncertainty estimation. In this subsection, we compare the proposed DE baseline against our model.
In the Tables 8 and 9 below, we present the results of this comparison. While the DE baseline delivers good accuracy and uncertainty quality, it is approximately 14 times slower than the single model baseline and about 5 times slower than our approach. These results confirm that while DE applied to the iterative architecture provides high-quality uncertainty estimates and accuracy, the computational cost is significantly higher compared to our approach. Our method strikes a balance between computational efficiency and performance, making it a more practical choice for real-world applications where computational resources and time are critical factors.
MC-DP Ens. Ens. Iter. Ours
Corr 87.1 87.4 88.1 85.2
Comp 58.2 66.7 76.1 77.8
Qual 54.1 60.8 69.1 68.6
F1 20.4 22.1 24.4 24.5
APLS 58.78 68.81 78.40 77.21
rAULC 30.18 72.19 69.03 69.23
Corr (unc.) 59.72 79.42 79.17 74.73
ECE 0.997 1.138 0.850 0.475
Train 1x 5x 14x 2.8x
Inf 5x 5x 13.5x 2.7x
Table 8: Uncertainty and performance metrics on RT Dataset. The best result in each category is in bold and the second best is in bold. Most correspond to Ours and Iterative Ensembles.
MC-DP Ens. Ens. Iter. Ours
Corr 81.6 83.6 93.9 92.3
Comp 92.3 90.4 85.9 86.7
Qual 78.2 78.7 81.3 81.1
F1 13.6 14.1 14.8 15.4
APLS 59.65 67.53 78.81 78.04
rAULC 19.56 78.65 69.03 79.27
Corr (unc.) 32.50 76.39 76.51 87.46
ECE 0.558 0.794 0.746 0.419
ROC-AUC 61.25 67.03 68.42 67.09
PR-AUC 62.64 67.85 71.25 72.11
Train 1x 5x 14x 2.8x
Inf 5x 5x 13.5x 2.7x
Table 9: Uncertainty and performance metrics on MS Dataset. The best result in each category is in bold and the second best is in bold. Most correspond to Ours and Iterative Ensembles.